OpenStudy (anonymous):

Can anyone help me think about a way to prove or disprove this? If f(x) >=0 and b >=0, then b*(integral from b to infinity of f(x)dx) <= (integral from b to infinity of x*f(x)dx).

OpenStudy (anonymous):

how about f(x)= e^-x b=1?

OpenStudy (anonymous):

nope that does not work

OpenStudy (anonymous):

seems rather unlikely to me

OpenStudy (anonymous):

try \(f(x)=\frac{1}{x^2}\)

OpenStudy (anonymous):

\[ \large b\int_b^\infty f(x)dx \leq \int_b^\infty xf(x)dx \]This?

OpenStudy (anonymous):

ya thats right wio

OpenStudy (anonymous):

then \[\int_b^\infty\frac{dx}{x^2}\] is finite, whereas \[\int_b^{\infty}\frac{dx}{x}\] is not

OpenStudy (anonymous):

oh i forgot the \(b\) out front no matter, take \(b=1\)

OpenStudy (anonymous):

right i saw that and was not sure if that would count as an answer. I then tried f(x) = 1/x^3

OpenStudy (anonymous):

all you need is one example to disprove, right?

OpenStudy (anonymous):

ok so because the first one is finite while the other one is infinite, that disproves the inequality?

OpenStudy (anonymous):

Since \(b\) is positive: \[ \large \int_b^\infty f(x)dx \leq \int_b^\infty \frac{x}{b}f(x)dx \]

OpenStudy (anonymous):

Maybe then try integration by parts to see if you can get something interesting.

OpenStudy (anonymous):

yes one is finite and the other is infinite that certainly disproves the inequality

OpenStudy (anonymous):

alright cool! Thank you i really appreciate it

OpenStudy (anonymous):

@satellite73 Really? So \(6 < \infty\) for example, is false... can be used as a counter example?

OpenStudy (anonymous):

actually it is a rather silly inequality, if it is what i read it to be if you take \(b=1\) then why on earth would \[\int_1^{\infty}f(x)dx\] be less than \[\int_1^{\infty}xf(x)dx\]? try with \(f(x)=\frac{1}{x^3}\) or anything of that type

OpenStudy (anonymous):

oh damn hold the phone i am way way wrong!!!

OpenStudy (anonymous):

ignore everything i wrote everything

OpenStudy (anonymous):

To me it looked like is was pretty right on.

OpenStudy (anonymous):

no i had the whole thing backwards. now i have to think maybe it is true

OpenStudy (anonymous):

after all, as \(x\to \infty\) at some point \(x>b\)

OpenStudy (anonymous):

Suppose \(c = x-b\) Then we have \(x = b+c\).

OpenStudy (anonymous):

\[ \large b\int_b^\infty f(x)dx \leq \int_b^\infty (b+c)f(x)dx \] This gets us: \[ \large b\int_b^\infty f(x)dx \leq b\int_b^\infty f(x)dx +c\int_b^\infty f(x)dx \]Then: \[\large 0 \leq c\int_b^\infty f(x)dx \]

OpenStudy (anonymous):

oops, let me redo that

OpenStudy (anonymous):

\[ \large b\int_b^\infty f(x)dx \leq b\int_b^\infty f(x)dx +\int_b^\infty (x-b) f(x)dx \]

OpenStudy (anonymous):

Since we start integrating at \(b\), it's clear that \(x-b\) is positive and \(f(x)\) is positive.

OpenStudy (anonymous):

So we could say \[ \large 0 \leq \int_b^\infty (x-b)f(x)dx \]

OpenStudy (anonymous):

why can we say that x-b is positive ?

OpenStudy (anonymous):

perhaps I should be saying 'non negative' rather than positive.

OpenStudy (anonymous):

Remember that an integral is a infinite sum. The sum is starting with \(x=b\) and ending with \(x=\infty\).

OpenStudy (anonymous):

ahhhh right ok

OpenStudy (anonymous):

You'll probably want to convert the integral to summation notation to show this.

OpenStudy (anonymous):

\[ \large \int_b^\infty (x-b)f(x)dx = \lim_{n \to \infty} \sum_i^n (x_i^*-b)f(x_i^*)\Delta x \]

OpenStudy (anonymous):

\[ b \leq x_i^* < \infty \]

OpenStudy (anonymous):

\[ 0 \leq x_i^* -b < \infty \]

OpenStudy (anonymous):

ok i think the summation notation is making it a little harder for me to understand

OpenStudy (anonymous):

i understand the whole integral part from above. is that good to show that it is true?

OpenStudy (anonymous):

it seems like it should be but im kind of new to proofs

OpenStudy (anonymous):

Ummm, I think it would be a bit iffy, but I think it should be okay.

OpenStudy (anonymous):

It's a matter of rigor. Perhaps I'm being too rigorous.

OpenStudy (anonymous):

alright i know what you mean

OpenStudy (anonymous):

its starting to make sense

OpenStudy (anonymous):

Basically I'm showing that you're adding a bunch of non-negative numbers, resulting in a non-negative number.

OpenStudy (anonymous):

Basically this proof will be doing everything we did here backwards.

OpenStudy (anonymous):

and because they can not be negative then if you multiply by x there is no way it will be less then if you dont multiply by x right?

OpenStudy (anonymous):

You want to start by asserting this: \[ \large 0 \leq \int_b^\infty (x-b)f(x)dx \]THEN you add \[ \large b\int_b^\infty f(x)dx \] To both sides.

OpenStudy (anonymous):

So that first assertion is key to proving it.